6x-9-2(1+x)=x-9
 
6x-9-2(1+x)=x-9
 
6x-9-2(1+x)=x-9
 
6x-9-2(1+x)=x-9

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

2(x+2)+5(x+5) = 4(x-8)+2(x-2)

2x + 4 + 5x + 25 = 4x – 32 + 2x – 4

2x + 5x -4x – 2x = -32 – 4 -4 – 25

7x – 6x = -65

x = – 65

Thanks